# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

### v1

@@ 3.3.2. Bidirectional Knowledge Distillation (BKD)
(5) $L_{di}^{KL}=KL(z_{di}\circ(x)\parallel\hat{z}_{dw}\circ(x))$
(6) $L_{dw}^{KL}=KL(\hat{z}_{dw}\circ(x)\parallel z_{di}\circ(x))$
@@
(7) $L=L^{CE}+\alpha L^{KL}$
3.3.3. Optimization process
As depicted in Algorithm 1, the updating process occurs in two phases: server-side and client-side. On the server side, the global domain-invariant adapter $A_{di}^{0}$ and the domain-aware adapters $A_{dw}^{k,0}$ for each client are initialized along with the initial weights $a_{k}^{0}$. In each communication round, if it is not the initial round, the server collects the generalization gap $G_{k}^{r}$ and updated domain-invariant adapters $A_{di}^{k,r}$ from all clients. The server then adjusts the weights $a_{k}^{r}$ based on client performance and aggregates this information using the GA strategy(Zhang et al., 2023c), represented as:
(8) $A_{\mathrm{inv}}^{(r+1)}=\sum_{k=1}^{K}\frac{G_{k}^{(r)}}{\sum_{j=1}^{K}G_{j}^{(r)}}\hat{A}_{\mathrm{inv},k}^{(r)},$
This iterative process of adapter refinement and knowledge distillation culminates in the broadcast of the updated global domain-invariant adapter to all clients, completing a synergistic cycle of domain-invariant feature consolidation and source-specific knowledge preservation. The continuous interplay between these dual adapters across communication rounds progressively enhances the model’s capacity to disentangle cross-domain patterns while retaining discriminative source characteristics.
4. Experimental Results
@@
4.1. Datsets
We evaluate the effectiveness of our method using four mainstream domain generalization datasets. Below is a description of each dataset.
PACS Dataset(Li et al., 2017): The PACS dataset consists of 9,991 images across seven object categories distributed among four visually distinct domains: Photo, Art Painting, Cartoon, and Sketch.
OfficeHome Dataset(Venkateswara et al., 2017): The OfficeHome dataset comprises 24,788 images spanning 65 categories across four domains: Art, Clipart, Product, and Real World.
VLCS Dataset(Fang et al., 2013): The VLCS dataset integrates images from four photographic collections: PASCAL VOC2007, LabelMe, Caltech-101, and SUN09, containing a total of 10,729 images across five common categories. .
DomainNet Dataset(Peng et al., 2019): DomainNet is the largest and most challenging benchmark for domain generalization, comprising 569,010 images across 345 categories distributed over six domains: Clipart, Infograph, Painting, Quickdraw, Real World, and Sketch.
For each of the datasets mentioned above, we employed a leave-one-domain-out testing strategy. Specifically, one domain is selected as the target domain for testing while the remaining domains serve as source domains for model training. Finally, we report the classification accuracy and overall average accuracy for each domain.
4.2. Baseline Methods
@@
We comprehensively compared our FedSDAF approach against three main categories of methods, encompassing a total of 14 baseline approaches:
(1) Centralized learning based Domain Generalization (DG) Methods. These methods operate by centrally collecting data from all available domains to train models with robust generalization capabilities. Specifically, we included the following approaches in this category: SWAD(Cha et al., 2021), which identifies flat minima through systematic random weight averaging, thereby stabilizing the learning process across different domains and enhancing generalization performance; HCVP(Zhou et al., 2025), which employs a hierarchical cue generation system that effectively integrates domain and task features, consequently enhancing cross-domain performance; and Doprompt(Zheng et al., 2022), which generates adaptive cues specifically tailored to target images, thus significantly improving generalization capabilities when applied to previously unseen domains.
(2) CNN-based Federated Domain Generalization (FedDG) Methods. These methods are based on the CNN architecture specifically developed to address the data segregation challenge in joint learning scenarios. We evaluated: FedSR(Nguyen et al., 2022), which learns generalized representations by effectively leveraging L2 regularity and mutual information constraints among federated clients, thereby promoting knowledge sharing while respecting data privacy; FedADG(Zhang et al., 2023b), which utilizes adversarial learning techniques to align source domains with reference distributions in a federated context, consequently enhancing domain generalization; CCST(Chen et al., 2023), which implements cross-client style migration to enable sharing of domain-specific knowledge between clients without directly exchanging raw data; ELCFS(Liu et al., 2021a), which facilitates information exchange between clients through frequency domain transformations, thus preserving privacy while enabling effective knowledge transfer; and GA(Zhang et al., 2023c), which incorporates a dynamic weight adjustment module on the server side to optimize the global model’s generalization capability across diverse domains.
(3) Parameter Efficient Fine-Tuning (PEFT) Methods. These methods integrate adapters for efficient fine-tuning, primarily aimed at improving model generalization in visual-language models (VLMs). Our comparison included: FedCLIP(Lu et al., 2023), which employs a CLIP-based attention adapter that effectively leverages visual-language alignment within a federated context; PromptFL(Guo et al., 2024), which learns client-independent textual cues after aggregation, enabling more efficient knowledge sharing while maintaining model performance; FedAPT(Su et al., 2024c), which collaboratively trains meta-cues and adaptive networks to significantly enhance generalization capabilities across heterogeneous domains; FedPR(Feng et al., 2023), which strategically optimizes local visual cues within a global cue zero space, achieving an optimal balance between local adaptation and global consistency; FedMaPLe(Khattak et al., 2023), which implements multimodal cue learning in a federated environment to effectively leverage diverse information sources; and PLAN(Gong et al., 2024), an innovative framework that utilizes reference-based prompt learning to synchronize local prompts across clients while selectively aggregating them into global prompts, thereby substantially enhancing generalization in federated domain generalization scenarios.
Table 2. Performance Comparison on PACS, VLCS, and OfficeHome Datasets (%)
Methods PACS Avg. VLCS Avg. OfficeHome Avg.
A C P S C L V S A C P R
@@
(PEFT) Methods(backbone: ViT-B/16)
@@
Note: Bold numbers indicate the best performance in each category. All values are in percentage (%).
Table 3. Performance Comparison on DomainNet Datasets (%)
@@ 4.4. Main Results
In our experimental evaluations, we achieved state-of-the-art (SOTA) results across all four mainstream datasets: PACS, VLCS, OfficeHome, and DomainNet. This accomplishment underscores the robust generalization capabilities of our proposed Federated Source Domain Awareness Framework (FedSDAF). Notably, our method demonstrated significant advancements on the S-domain of the PACS dataset, where previous approaches had struggled to deliver satisfactory outcomes. Specifically, our framework achieved an impressive accuracy of 98.62%, reflecting a notable improvement of 3.6
@@
Existing federated domain generalization (FedDG) methods, such as FedSR and FedADG, exhibit limitations, particularly in their reliance on local data without sufficient integration of domain-specific knowledge from diverse sources. While these methods have made strides in representation learning, they often fall short in achieving robust generalization across unseen domains. In contrast, our framework incorporates the utilization of source-domain perceptual features, which have proven instrumental in enhancing model adaptability to unseen target domains. This is particularly evident in our results, where the integration of domain-aware features has led to significant performance gains. Additionally, our bidirectional knowledge distillation mechanism facilitates effective knowledge sharing between clients, thereby enhancing model adaptability while preserving privacy.
4.5. Ablation Experiment
Table 4. TABLE VI INVESTIGATION ON THE CONTRIBUTION OF DIFFERENT COMPONENTS
@@ 4.5. Ablation Experiment
4.6. Convergence Analysis
Figure 4. Convergence analysis on different clients in Domain benchmark
@@ 4.6. Convergence Analysis
A woman and a girl in white dresses sit in an open car.
To validate the convergence properties of our FedSDAF framework, we conduct a comparative analysis against the baseline FedAVG (Adapter) method using the PACS dataset, which encompasses Art, Photo, Cartoon, and Sketch domains. As demonstrated in Figure 4, FedSDAF exhibits superior convergence characteristics in terms of both speed and stability across all domains, particularly during the initial communication rounds. FedSDAF demonstrates substantial early-stage convergence, achieving accuracy improvements of 5.8-7.2% over FedAVG within the first 10 communication rounds. This accelerated convergence can be attributed to the domain-aware adapter’s MHSA mechanism, which effectively prioritizes domain-specific feature extraction during local client updates. In contrast, FedAVG’s conventional adapter aggregation exhibits slower convergence, requiring additional rounds to reconcile heterogeneous domain representations due to its domain-agnostic nature. The convergence analysis extends to Round 40, where FedSDAF demonstrates stable convergence across all PACS domains with accuracy rates ranging from 93.6% to 97.1%, consistently outperforming FedAVG by margins of 9.4-13.8%. This enhanced convergence stability is primarily achieved through the bidirectional knowledge distillation mechanism, which facilitates systematic feature transfer between domains. Specifically, the framework’s ability to integrate diverse domain characteristics, such as texture-rich features from the Art domain with shape-centric patterns from the Sketch domain, contributes to its faster and more stable convergence trajectory. Unlike FedAVG’s static aggregation strategy, which often struggles to achieve stable convergence due to domain heterogeneity, FedSDAF’s iterative distillation process ensures consistent convergence behavior while effectively leveraging cross-domain complementarity.
4.7. Hyperparameter Analysis
Figure 5. Effects of $\alpha$.
4.8. Visualization
Figure 6. Visualization of the target features extracted by ViT and FedSDAF.
@@ 4.8. Visualization
A woman and a girl in white dresses sit in an open car.

### v3

@@ 3.3.2 Bidirectional Knowledge Distillation (BKD)
$L_{di}^{KL}=KL(z_{di}\circ(x)\parallel\hat{z}_{dw}\circ(x))$ (3)
$L_{dw}^{KL}=KL(\hat{z}_{dw}\circ(x)\parallel z_{di}\circ(x))$ (4)
@@
$L=L^{CE}+\alpha(L_{di}^{KL}+L_{dw}^{KL})$ (5)
3.3.3 Optimization process
The optimization process unfolds across the server and clients. On the server, the global domain-invariant adapter $A_{di}$ is initialized. In each communication round $r$, the server distributes the current global adapter $A_{di}^{r}$ to a subset of clients. Each participating client $k$ receives $A_{di}^{r}$ and performs local training on its dataset $\mathcal{D}_{k}$. During local training, the client updates its local copy of the domain-invariant adapter, $\hat{A}_{di}^{k,r}$, and its private domain-aware adapter, $A_{dw}^{k}$, using the bidirectional knowledge distillation loss described in Equation (5). After local training, each client sends its updated adapter $\hat{A}_{di}^{k,r}$ back to the server. The server then aggregates the received adapters to produce the next round’s global adapter, $A_{di}^{r+1}$, using standard federated averaging:
$A_{di}^{r+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{N}\hat{A}_{di}^{k,r}$ (6)
where $n_{k}$ is the number of samples on client $k$ and $N=\sum_{k}n_{k}$ is the total number of samples across all clients. This iterative process allows the framework to build a robust global consensus ($A_{di}$) while leveraging the rich, localized knowledge preserved within each client’s $A_{dw}$.
4 Experimental Results
@@
4.1 Datasets and Baselines
Datasets. We evaluate our method on four standard DG benchmarks: PACS [21], OfficeHome [22], VLCS [23], and the large-scale DomainNet [24]. For all datasets, we employ a leave-one-domain-out evaluation protocol, where one domain is held out as the unseen target for testing.
@@ 4.1 Datasets and Baselines
Baseline Methods. We compare FedSDAF against 14 methods from three categories:
  1. 1.
    Centralized DG Methods: These methods require centralized data access. Baselines include SWAD [25], HCVP [26], and Doprompt [27].
  2. 2.
    CNN-based FedDG Methods: These methods use CNN backbones in a federated setting. Baselines include FedSR [28], FedADG [29], CCST [11], ELCFS [9], and GA [20].
  3. 3.
    Parameter-Efficient Fine-Tuning (PEFT) Methods: These recent methods are most comparable to our approach. Baselines include FedCLIP [30], PromptFL [31], FedAPT [32], FedPR [33], FedMaPLe [34], and PLAN [53].
Table 2: Performance Comparison on PACS, VLCS, and OfficeHome Datasets (%)
Methods PACS VLCS OfficeHome
A C P S Avg. C L V S Avg. A C P R Avg.
@@
PEFT Methods (backbone: ViT-B/16)
@@
Table 3: Performance Comparison on DomainNet Datasets (%)
@@ 4.3 Main Results
In our experimental evaluations, we achieved state-of-the-art (SOTA) results across all four mainstream datasets: PACS, VLCS, OfficeHome, and DomainNet. This accomplishment underscores the robust generalization capabilities of our proposed Federated Source Domain Awareness Framework (FedSDAF). Notably, our method demonstrated significant advancements on the S-domain of the PACS dataset, where previous approaches had struggled to deliver satisfactory outcomes. Specifically, our framework achieved an impressive accuracy of 98.62%, reflecting a notable improvement of 3.6% over prior methods.
@@
Existing federated domain generalization (FedDG) methods, such as FedSR and FedADG, exhibit limitations, particularly in their reliance on local data without sufficient integration of domain-specific knowledge from diverse sources. While these methods have made strides in representation learning, they often fall short in achieving robust generalization across unseen domains. In contrast, our framework incorporates the utilization of source-domain perceptual features, which have proven instrumental in enhancing model adaptability to unseen target domains. This is particularly evident in our results, where the integration of domain-aware features has led to significant performance gains. Additionally, our bidirectional knowledge distillation mechanism facilitates effective knowledge sharing between clients, thereby enhancing model adaptability.
4.4 Ablation Experiment
Table 4: Ablation Study for Different Components
@@ 4.4 Ablation Experiment
4.5 Convergence Analysis
Figure 3: Convergence analysis on different clients in Domain benchmark.
To validate the convergence of our FedSDAF framework, we compared it against the FedAVG (Adapter) baseline on the PACS dataset. As shown in Figure 3, FedSDAF converges significantly faster and more stably. Within the first 10 rounds, FedSDAF already achieves a 5.8-7.2% accuracy advantage. This rapid start is driven by the local domain-aware adapter’s MHSA mechanism, which accelerates the extraction of domain-specific features, whereas FedAVG’s domain-agnostic aggregation lags behind. By round 40, FedSDAF’s lead widens to a stable 9.4-13.8% accuracy margin. This superior stability is a direct result of our bidirectional knowledge distillation, which systematically fuses complementary knowledge across domains. This dynamic process overcomes the limitations of static aggregation strategies that are typically hindered by domain heterogeneity, ensuring a consistently faster and more robust convergence.
4.6 Hyperparameter Analysis
Figure 4: Effects of $\alpha$.
4.7 Communication Cost Analysis
A concern regarding our framework could be the complexity introduced by the MHSA mechanism. However, this design choice does not translate to increased communication overhead. The Domain-Aware Adapter, which contains the MHSA module, is kept strictly local to each client and is never aggregated by the server. Therefore, its parameters do not contribute to the communication cost.
The communication cost per round is dominated by two components: (1) the upload/download of the lightweight Domain-Invariant Adapter ($\hat{A}_{di}$), and (2) the exchange of logits for the Bidirectional Knowledge Distillation (BKD) process. As shown in Figure 5, the size of the logits is negligible (0.003 MB). The total estimated communication cost for FedSDAF per client per round is approximately 2.270 MB. This is not only comparable to other PEFT-based methods like FedClip (2.010 MB) but is also an order of magnitude smaller than full-model aggregation methods like FedAvg (13.650 MB). This analysis demonstrates that FedSDAF remains highly communication-efficient, highlighting its practical potential for deployment in real-world federated learning scenarios.
Figure 5: Communication Cost Comparison per Client. FedSDAF’s total cost is comparable to other adapter-based methods and significantly lower than full-model federation.
5 Conclusion
In this paper, we introduced FedSDAF, a novel framework for Federated Domain Generalization that challenges the conventional focus on domain invariance. Our work is motivated by the key insight, validated through a targeted study, that source-domain-aware features possess superior generalization power in data-isolated environments. FedSDAF operationalizes this insight through an innovative dual-adapter architecture, which decouples and synergizes client-specific "local expertise" and a shared "global consensus." A Bidirectional Knowledge Distillation mechanism enables a collaborative dialogue between these components, allowing the framework to systematically transform unique source domain knowledge into a powerful signal for generalization. Extensive experiments on four challenging benchmarks demonstrate that FedSDAF significantly outperforms existing methods, establishing a new state-of-the-art.
Despite its strong performance, our approach has limitations. The dual-adapter architecture and bidirectional distillation process introduce a moderate increase in computational overhead on the client side during local training compared to simpler fine-tuning methods. Furthermore, the performance of FedSDAF, like other PEFT-based approaches, is inherently tied to the quality of the pre-trained foundation model. Future research could explore more computationally efficient distillation techniques or investigate the applicability of the source-aware paradigm to other modalities and federated learning challenges.
Appendix A Algorithm Details
As depicted in Algorithm 1, on the client side, each client initializes its local copy $\hat{A}_{di}^{k,r}$ corresponding to the received global domain-invariant adapter $A_{di}^{r}$. The client subsequently optimizes its local domain-aware adapter $A_{dw}^{k,r}$ and the local replica $\hat{A}_{di}^{k,r}$ through the bidirectional knowledge distillation mechanism. During this optimization, the client computes its training error $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$ and, after each round, uploads the generalization gap $G_{k}^{r}$ (if $r>0$) along with the updated domain-invariant adapters $A_{di}^{k,r}$ back to the server. This systematic updating process enables the FedSDAF method to effectively leverage source domain knowledge, thereby enhancing the model’s generalization performance in unseen target domains.
Algorithm 1 Federated Domain Generation with Bidirectional Distillation (FedSDAF)
1:$K$: Number of clients; $\{D_{k}\}_{k=1}^{K}$: Client datasets; $R$: Communication rounds
@@ Appendix A Algorithm Details
3:
4:Server Execution:
5:Initialize $A_{di}^{0}$, $A_{dw}^{k,0}\leftarrow$ random initialization
6:Collect $\{n_{k}=|D_{k}|\}_{k=1}^{K}$ from clients
@@
[…]
9:for $r\leftarrow 0$ to R1R-1 do
10:  if $r>0$ then
11:   Collect $\{A_{di}^{k,r}\}_{k=1}^{K}$
12:   $A_{di}^{r+1}\leftarrow\sum_{k=1}^{K}w_{k}\cdot A_{di}^{k,r}$
13:  else
14:   $A_{di}^{r+1}\leftarrow A_{di}^{0}$
15:  end if
16:  Broadcast $A_{di}^{r+1}$
17:end for
18:
19:Client $k$ Execution (Parallel):
20:Send $n_{k}=|D_{k}|$ to server
21:for $r\leftarrow 0$ to R1R-1 do
22:  Receive $A_{di}^{r}$
23:  $\hat{A}_{di}^{k,r}\leftarrow A_{di}^{r}$
24:  while not converged do
25:   $\min_{A_{dw}^{k,r},\hat{A}_{di}^{k,r}}\mathcal{L}_{\text{distill}}(A_{dw}^{k,r},\hat{A}_{di}^{k,r})$
26:  end while
27:  Send $A_{di}^{k,r}$
28:end for
Appendix B Supplementary Experiment
B.1 Fine-grained Ablation Study on Hard Domains
A potential concern is that the performance gain from components like MHSA appears modest when averaged across all domains. This is because performance on many domains is already saturated (e.g., >99% accuracy), masking the components’ true impact. To provide a clearer picture, we conduct a fine-grained ablation study focusing on the most challenging domains from three benchmarks: PACS (Photo), VLCS (LabelMe), and OfficeHome (Clipart), where models typically struggle the most.
The results are presented in Table 5. On these hard domains, the contribution of each component becomes significantly more pronounced. For instance, enabling BKD and MHSA leads to substantial gains of 3-7% on VLCS and OfficeHome. The full model consistently achieves the best results, with the MHSA mechanism being particularly crucial for boosting performance on these challenging generalization tasks. This underscores the importance of each component in our framework for achieving robust generalization, an effect that is most visible where it matters most.
Table 5: Fine-grained ablation study on the most challenging domains. The performance improvements for each component are more significant on these domains compared to the overall average.
Components Accuracy (%) on Hard Domains
BKD MHSA $\boldsymbol{\hat{A}_{di}}$ $\boldsymbol{\hat{A}_{dw}}$ PACS (P) VLCS (L) OfficeHome (C)
99.28 85.09 76.20
$\times$ 98.53 82.15 73.45
$\times$ 98.21 81.88 72.93
$\times$ 97.55 79.50 70.10
$\times$ 97.98 80.20 71.30
$\times$ $\times$ $\times$ $\times$ 95.10 74.30 68.50
B.2 Visualization
Figure 6: Visualization of the target features extracted by ViT and FedSDAF.